# The Global Sections of Chiral de Rham Complexes on Closed Complex Curves

### v1

@@ 1. Introduction
The space of global sections $\Gamma(X,\Omega_{X}^{ch})$ is always a vertex superalgebra, and it is known to have extra structure when $X$ is endowed with geometric structures. For example, when $X$ is Kähler or hyperkähler, it has an $\mathcal{N}=2$ structure or $\mathcal{N}=4$ structure, respectively [2]. Recently in [17, 11], Linshaw and the first auther gave a complete description of $\Gamma(X,\Omega_{X}^{ch})$ for any compact Ricci-flat Kähler manifold $X$. In [4], for any congruence subgroup $G\subset SL(2,\mathbb{R})$, Dai constructed a basis of the $G$-invariant global sections of the chiral de Rham complex on the upper half plane, which are holomorphic at the cusps. The vertex operations are determined by a modification of the Rankin-Cohen brackets of modular forms. Further in [5], it is shown that if the global sections are allowed to be meromorphic at cusp, the space of $G$-invariant global sections will be a simple vertex algebra. In this paper, we extend the method developed in [17] to closed local symmetric space and calculate the space of global sections of chiral de Rham complex on any closed curves with genus $g\geq 2$ (Theorem 9).
@@ 2.1. Vertex algebras
Using the above construction, we abtain a smooth version of the chiral de Rham complex $\Omega_{X}^{ch,sm}$ on $X$, by regarding $X$ as a smooth real manifold, and replacing $\mathcal{W}(\mathbb{C}^{d})$, the complex coordinate system and the space of analytic functions $\mathcal{O}(U)$ by $\mathcal{W}(\mathbb{C}^{2d})$, the smooth coordinate system and the space of smooth complex functions $C^{\infty}(U)$, respectively.
@@ 3. Antiholomorphic vector bundles corresponding to chiral de Rham complexes
Let $\theta_{ij}=-\sum_{k}H^{ik}\bar{\partial}H_{kj}$ be a (0, 1)-form, then:
@@
For any open subset $U$ of $X$, construct a homomorphism $I_{U}$ locally:
@@
for any $a\in\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))(U)$. According to the above Theorem 2 and Theorem 1, we get the following theorem,
Theorem 3.
$H^{i}(X,\Omega^{ch,*}_{X})\cong H^{i}(\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))(X),\bar{D}).$
@@
$\Gamma(X,\Omega^{ch}_{X})=H^{0}(X,\Omega^{ch,*}_{X})\cong H^{0}(\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))(X),\bar{D}).$
@@ 3.4. Hermitian locally symmetric space
$R(\cdot,\cdot)\frac{\partial}{\partial{\gamma^{j}}}=\bar{\partial}(\partial H_{jk}H^{ki})\frac{\partial}{\partial{\gamma^{i}}}=-\bar{\partial}\bar{\theta}_{ij}\frac{\partial}{\partial{\gamma^{i}}}=-\overline{\partial\theta_{ij}}\frac{\partial}{\partial{\gamma^{i}}}.$
@@
$\overline{\partial(\mathbf{B}^{k}_{(0)}\theta_{ij})}\frac{\partial}{\partial{\gamma^{i}}}=\bar{\partial}(\overline{\mathbf{B}^{k}_{(0)}\theta_{ij}}\frac{\partial}{\partial{\gamma^{i}}})=-\bar{\partial}(R(\cdot,\overline{H^{kl}\frac{\partial}{\partial{\gamma^{l}}}})\frac{\partial}{\partial{\gamma^{j}}})=-(\nabla_{\bar{\partial}}R)(\cdot,\overline{H^{kl}\frac{\partial}{\partial{\gamma^{l}}}})\frac{\partial}{\partial{\gamma^{j}}}=0.$
@@
$\theta=-H^{-1}\bar{\partial}H=-\frac{2}{\gamma-\bar{\gamma}}d\bar{\gamma},\quad\mathbf{B}_{(0)}\theta=id\bar{\gamma}.$
Thus by Lemma 1, we have
Corollary 1.
@@
Here $F_{2}=\tilde{F}_{2}H^{-1}\nabla_{\frac{\partial}{\partial{\gamma}}}$ and $\tilde{F}_{2}$ commutes with $\nabla_{\frac{\partial}{\partial\gamma}}$ as in the proof of Lemma 3. ∎
@@ 4. Global Sections on closed Complex Curves
In this section we use Theorem 3 to calculate the space of global sections of chiral de Rham complex on closed curves with genus $g\geq 2$.
@@
where $a_{s-i}$ is a smooth section of $SW(\bar{T}^{*}X)[k,l,s-i]$ and $a_{s}$ is nonzero. The summation is finite since $SW(\bar{T}^{*}X)[k,l,s-i]=\mathbf{0}$ when $k<i-s$. By Corollary 1 , $\bar{D}a=0$ is divided into a series of equations bellow:
@@
When $l-s<0$, by Lemma 5, the mean curvature of $SW(\bar{T}^{*}X)[k,l,s]$ is negative-definite everywhere on $X$, so by above Theorem 4, there is no nonzero holomorphic section $a_{s}$.
@@
The last equality follows from Lemma 2 and the fact that $a_{s}$ is a holomorphic section. So $a_{s-1}=\frac{-1}{l-s}\bar{\partial}^{\prime*}F_{1}a_{s}$ is a solution of the equation. ∎
Lemma 7.
Suppose $a_{s-1}$ is given as in Lemma 6, then $a_{s-2}=\frac{-1}{2l-2s+1}(\bar{\partial}^{\prime*}F_{1}a_{s-1}+\bar{\partial}^{\prime*}F_{2}a_{s})$ is a solution of the equation ¯$\bar{\partial}^{\prime}a_{s-2}+F_{1}a_{s-1}+F_{2}a_{s}=0$.
@@
Using the expression in Lemma 6, we obtain:
@@
For the term ¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s}$, using Lemma 4, we have:
@@
Induction on $t$. The cases $t=1,2$ are verified in Lemma 6 and Lemma 7, respectively. Next assume that $t\geq 3$.
@@
This follows directly from equation (3.10) and Lemma 8. ∎
@@
As in the proof of Lemma 6, we have:
@@
That is $F_{1}a_{s}=0$ by Theorem 6. ∎
Theorem 7.
@@
By Lemma 9, if there is an $a=\sum_{i\geq 0}a_{s-i}$ with $\bar{D}a=0$, we must have $F_{1}a_{s}=0$. By Corollary 1 and Corollary 2, if $F_{1}a_{s}=0$, then $\bar{D}a_{s}=0$. ∎
@@ 4.3. The $\mathfrak{sl}_{2}$ action on a $\beta\gamma-bc$ system
By Lemma 8, the space of holomorphic sections of $\bigoplus_{k,s}SW(\bar{T}^{*}X)[k,s,s]$ is isomorphic to its fibre $\bigoplus_{k,s}SW^{\gamma}[k,s,s]$. That is the subspace of $SW^{\gamma}$ generated by monomials with the number of $\mathbf{B}$’s plus the number of $\mathbf{b}$’s is equal to the number of $\mathbf{\Gamma}$’s plus the number of $\mathbf{c}$’s. By Lemma 10,

### v2

@@ 1. Introduction
The space of global sections $\Gamma(X,\Omega_{X}^{ch})$ is always a vertex superalgebra, and it is known to have extra structure when $X$ is endowed with geometric structures. For example, when $X$ is Kähler or hyperkähler, it has an $\mathcal{N}=2$ structure or $\mathcal{N}=4$ structure, respectively [2]. Recently in [17, 11], Linshaw and the first auther gave a complete description of $\Gamma(X,\Omega_{X}^{ch})$ for any compact Ricci-flat Kähler manifold $X$. In [4], for any congruence subgroup $G\subset SL(2,\mathbb{R})$, Dai constructed a basis of the $G$-invariant global sections of the chiral de Rham complex on the upper half plane, which are holomorphic at the cusps. The vertex operations are determined by a modification of the Rankin-Cohen brackets of modular forms. Further in [5], it is shown that if the global sections are allowed to be meromorphic at cusp, the space of $G$-invariant global sections will be a simple vertex algebra. In this paper, we extend the method developed in [17] to closed local symmetric space and calculate the space of global sections of chiral de Rham complex on any closed curves with genus $g\geq 2$ (Theorem 4.6).
@@ 2.1. Vertex algebras
Using the above construction, we obtain a smooth version of the chiral de Rham complex $\Omega_{X}^{ch,sm}$ on $X$, by regarding $X$ as a smooth real manifold, and replacing $\mathcal{W}(\mathbb{C}^{d})$, the complex coordinate system and the space of analytic functions $\mathcal{O}(U)$ by $\mathcal{W}(\mathbb{C}^{2d})$, the smooth coordinate system and the space of smooth complex functions $C^{\infty}(U)$, respectively.
@@ 3. Antiholomorphic vector bundles corresponding to chiral de Rham complexes
Let $\theta_{ij}=-\sum_{k}H^{ik}\bar{\partial}^{\prime}H_{kj}$ be a (0, 1)-form, then:
@@
For any open subset $U$ of $X$, construct a vector space homomorphism $I_{U}$ locally:
@@
for any $a\in\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))(U)$. According to the above Theorem 3.1 and Theorem 2.1, we get the following theorem,
Theorem 3.2.
$H^{i}(X,\Omega^{ch}_{X})\cong H^{i}(\Omega^{ch,*}_{X},\bar{\partial})\cong H^{i}(\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))(X),\bar{D}).$
@@
$\Gamma(X,\Omega^{ch}_{X})\cong H^{0}(\Omega^{ch,*}_{X},\bar{\partial})\cong H^{0}(\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))(X),\bar{D}).$
@@ 3.4. Hermitian locally symmetric space
$R(\cdot,\cdot)\frac{\partial}{\partial{\gamma^{j}}}=\bar{\partial}^{\prime}(\partial H_{jk}H^{ki})\frac{\partial}{\partial{\gamma^{i}}}=-\bar{\partial}^{\prime}\bar{\theta}_{ij}\frac{\partial}{\partial{\gamma^{i}}}=-\overline{\partial\theta_{ij}}\frac{\partial}{\partial{\gamma^{i}}}.$
@@
$\overline{\partial(\mathbf{B}^{k}_{(0)}\theta_{ij})}\frac{\partial}{\partial{\gamma^{i}}}=\bar{\partial}^{\prime}(\overline{\mathbf{B}^{k}_{(0)}\theta_{ij}}\frac{\partial}{\partial{\gamma^{i}}})=-\bar{\partial}^{\prime}(R(\cdot,\overline{H^{kl}\frac{\partial}{\partial{\gamma^{l}}}})\frac{\partial}{\partial{\gamma^{j}}})=-(\nabla_{\bar{\partial}^{\prime}}R)(\cdot,\overline{H^{kl}\frac{\partial}{\partial{\gamma^{l}}}})\frac{\partial}{\partial{\gamma^{j}}}=0.$
@@
$\theta=-H^{-1}\bar{\partial}^{\prime}H=-\frac{2}{\gamma-\bar{\gamma}}d\bar{\gamma},\quad\mathbf{B}_{(0)}\theta=id\bar{\gamma}.$
Thus by Lemma 3.1, we have
Corollary 3.1.
@@
Here $F_{2}=\tilde{F}_{2}H^{-1}\nabla_{\frac{\partial}{\partial{\gamma}}}$ and $\tilde{F}_{2}$ commutes with $\nabla_{\frac{\partial}{\partial\gamma}}$ as in the proof of Lemma 3.3. ∎
@@ 4. Global Sections on closed Complex Curves
In this section we use Theorem 3.2 to calculate the space of global sections of chiral de Rham complex on closed curves with genus $g\geq 2$.
@@
where $a_{s-i}$ is a smooth section of $SW(\bar{T}^{*}X)[k,l,s-i]$ and $a_{s}$ is nonzero. The summation is finite since $SW(\bar{T}^{*}X)[k,l,s-i]=\mathbf{0}$ when $k<i-s$. By Corollary 3.1 , $\bar{D}a=0$ is divided into a series of equations bellow:
@@
When $l-s<0$, by Lemma 4.1, the mean curvature of $SW(\bar{T}^{*}X)[k,l,s]$ is negative-definite everywhere on $X$, so by above Theorem 4.1, there is no nonzero holomorphic section $a_{s}$.
@@
The last equality follows from Lemma 3.2 and the fact that $a_{s}$ is a holomorphic section. So $a_{s-1}=\frac{-1}{l-s}\bar{\partial}^{\prime*}F_{1}a_{s}$ is a solution of the equation. ∎
Lemma 4.3.
Suppose $a_{s-1}$ is given as in Lemma 4.2, then $a_{s-2}=\frac{-1}{2l-2s+1}(\bar{\partial}^{\prime*}F_{1}a_{s-1}+\bar{\partial}^{\prime*}F_{2}a_{s})$ is a solution of the equation ¯$\bar{\partial}^{\prime}a_{s-2}+F_{1}a_{s-1}+F_{2}a_{s}=0$.
@@
Using the expression in Lemma 4.2, we obtain:
@@
For the term ¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s}$, using Lemma 3.4, we have:
@@
Induction on $t$. The cases $t=1,2$ are verified in Lemma 4.2 and Lemma 4.3, respectively. Next assume that $t\geq 3$.
@@
This follows directly from equation (3.10) and Lemma 4.4. ∎
@@
As in the proof of Lemma 4.2, we have:
@@
That is $F_{1}a_{s}=0$ by Theorem 4.3. ∎
Theorem 4.4.
@@
By Lemma 4.5, if there is an $a=\sum_{i\geq 0}a_{s-i}$ with $\bar{D}a=0$, we must have $F_{1}a_{s}=0$. By Corollary 3.1 and Corollary 4.1, if $F_{1}a_{s}=0$, then $\bar{D}a_{s}=0$. ∎
@@ 4.3. The $\mathfrak{sl}_{2}$ action on a $\beta\gamma-bc$ system
By Lemma 4.4, the space of holomorphic sections of $\bigoplus_{k,s}SW(\bar{T}^{*}X)[k,s,s]$ is isomorphic to its fibre $\bigoplus_{k,s}SW^{\gamma}[k,s,s]$. That is the subspace of $SW^{\gamma}$ generated by monomials with the number of $\mathbf{B}$’s plus the number of $\mathbf{b}$’s is equal to the number of $\mathbf{\Gamma}$’s plus the number of $\mathbf{c}$’s. By Lemma 4.6,